12. Refer to the table below. According to the table, Julia has the absolute advantage in:

Time to make a pie Time to make a cake

Martha: 60 minutes 80 minutes

Julia: 50 minutes 60 minutes

A) pies B) both pies and cakes C) neither pies nor cakes D) cakes

13. Suppose it takes Dan 5 minutes to make a sandwich and 15 minutes to make a smoothie, and it takes Tracy 6 minutes to make a sandwich and 12 minutes to make a smoothie. What is the opportunity cost to Dan of making a sandwich?

A) 3 smoothies B) 15 smoothies C) 1/3 of a smoothie   D) 5 smoothies

14. In general, individuals and nations should specialize in producing those goods for which they have a(n):
A)   comparative advantage.
B)   absolute advantage.
C)   absolute advantage and a comparative advantage.
D)   absolutely comparative advantage.

15. A country is most likely to have a comparative advantage in the production of cars if:
A)   it has strict environmental protection laws governing automobile emissions.
B)   it has a relative abundance in the natural resources needed to produce cars.
C)   its citizens prefer driving cars to other forms of transportation.
D)   it imports most of the raw materials necessary to produce cars.

16. Smith and Jones comprise a two-person economy. Their hourly rates of production are shown below.

Calculation Per hour Computers per hour

Smith: 100 10

Jones: 120 6

A) consumption possibilities curve.   B) production function. C) supply curve. D) production possibilities curve.
Answer #1

12) Both pies and cakes

( Because she takes less time to bake a cake and pie compared to Martha. )

13) 1/3 of coffee

( Dan takes 5 minutes to make a sandwich and 15 minutes to make a smoothie.

Opportunity cost of making a sandwich = 5/15= 1/3 cup of smoothie )

14) comparative advantage

( It makes a nation able to specialise and trade even if it doesn't have absolute advantage in the production of both goods.)

15) lt has a relative abundance of natural resources needed to produce cars.

( It will enable the nation to produce more at less costs.)

16) production possibilities curve

( The given data shows production possibilities of the two person economy. )